OpenStudy (anonymous):

solve (x-1) ----- -1/4(x-2)=1 3

OpenStudy (ash2326):

Is this your question \[\frac{x-1}{3}-\frac{1}{4(x-2)}=1\]

OpenStudy (anonymous):

\[(x-1)/3-1(x-2)/4=1\] \[4(x-1)-3(x-2)/12=1\] \[4x-4-3x+6=12\] \[x+2=12\] \[x=12-2=10\]

OpenStudy (anonymous):

@Parvathy-Is (x-2) in denominator or numerator?